OpenStudy (chillout):

It's a quite long optimization question... Consider 2 straight lines: y=1 and y=-1, and the point A(0,3) in the xy-plane. Take a point P on the straight line y=1 and a point Q on the line y=-1 such that PÄQ = 90°. Let the two points P and Q move preservering the above conditions. We are to find the minimum value of lenght of the line PQ. First, denote the coordinates of P by (a,1) and Q by (b,-1). Then the condition PÄQ=90° is reduced to the conditions a!=0 and b!=0 and ab = (?, first question) Since we know that a and b have opposite signs, let us assume that a<0

OpenStudy (chillout):

Then we have PQ² = (b-a)² + (?, second question) = a²+b² + (?, third question) \[\ge2|ab| + 9(?) = (?)\]\[PQ \ge?\] Hence, when a = ? and b = ?, PQ takes the minimum value (?)

OpenStudy (chillout):

oh, and its 2|ab| + (?)

OpenStudy (jhannybean):

@Luigi0210 help him outt! :)

OpenStudy (anonymous):

what can be the condition for the required length to be minimum?

OpenStudy (chillout):

I've realised 1 thing so far. P is in the 2nd quadrant and Q is in the 4st quadrant. I don't know how this will help me, though. Well, it's just a guess but the length of one of the sides must be maximized. I'm pretty bad at optimization problems

OpenStudy (anonymous):

do you have the answer?

OpenStudy (chillout):

Yes, I do.

OpenStudy (anonymous):

is it \(\large 2\sqrt2 \) and the minimum length of PQ is 6 ???

OpenStudy (chillout):

Sec, gonna check.

OpenStudy (chillout):

Yes, the minimum value is 6

OpenStudy (anonymous):

cool... that wasnt so bad....

OpenStudy (chillout):

a is -2sqrt(2), b is 2sqrt(2) But I can't even start! gosh...

OpenStudy (anonymous):

right... ok....

OpenStudy (anonymous):

you'll need to find the relation between a and b... you can do this with the pythagorean theorem AND distance formula... just a sec for the drawing...

OpenStudy (anonymous):



OpenStudy (anonymous):

first, find: d1 = d2 = ???

OpenStudy (chillout):

Seems ok so far. I got here before but was stuck in the first question (a*b=?)

OpenStudy (anonymous):

that's what we're gonna get to.... the relation between a and b... after doing some substitutions, you should get \(\large a \cdot b=-8 \) or \(\large a=\frac{-8}{b} \)

OpenStudy (chillout):

well, d1: \[\sqrt{(0-a)² +(3-1)²}\] d2: \[\sqrt{(0-b)² +(3+1)²}\]

OpenStudy (anonymous):

ok good... simplified, we get: \(\large d_1=\sqrt{a^2+4} \) and \(\large d_2=\sqrt{b^2+16} \) correct?

OpenStudy (chillout):

Yes. That's it.